OpenStudy (anonymous):

Identify the vertex for the graph of y = −3x2 + 6x − 4.

OpenStudy (anonymous):

y = −3x2 + 6x − 4. y = −3x2 + 6x − 3-1 . y=-3(x^2-2x+1) -1 y+1=-3(x-1)^2 clearly the vertex is (1,-1)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

In these types of sums first try to complete the squares or the cube as the case may be and then use suitable transformation (translation of origin)

OpenStudy (anonymous):

yw